# Predicting and Understanding College Student Mental Health with Interpretable Machine Learning

### v1

@@ 5.2. Interpretability
Figure 6. The individual-level heatmaps that highlight feature importance for each interaction label across 217 UIDs. The y-axis represents label-specific features, which vary across different interaction labels.
@@ 5.2.1. Evaluation of Stage 1: Input Features $\rightarrow$ Interaction Labels
  • Label: Leisure - Feature F4 (Duration of walking) emerged as the most important predictor for 90.36% of UIDs, emphasizing walking as a key leisure activity. Other features, such as F13 (Phone usage at home), F3 (Duration of running), F5 (Duration of conversations), F8 (Voice detection at others’ dorms), and F14 (Phone usage at others’ dorms), are moderately important for 20–25% of UIDs, indicating that leisure also encompasses social interactions and phone use. In contrast, features like F1 (Biking duration), F2 (Footsteps), F6 (Duration of calls), F9 (Conversations at others’ dorms), F10 (Audio detections at social places), F11 (Duration at “leisure”), and F12 (Workout duration) consistently show lower relevance. F11 has low importance despite its direct link to the label, likely due to unclear dataset definitions, underscoring the need for improved data clarity. These findings illustrate how behaviors like walking and social interactions influence leisure and provide insights into broader themes related to mental health and lifestyle habits.
  • Label: Me Time - Features such as F16 (Phone usage at own dorm), F12 (Time spent at own dorm), F7 (Voice detections at own dorm), F13 (Study duration), and F17 (Phone usage at study locations) are important for 89.5% UIDs, with F12 and F16 being the most significant. These findings suggest that activities in dorm spaces, such as studying and socializing, are key indicators of “Me Time” and contribute to a supportive environment. Conversely, features like F1 (Biking duration), F2 (Footsteps), F3 (Running duration), F5 (Walking duration), and F14 (Workout duration) show low importance, indicating that individual physical activities are less relevant to this label. Other features display varying levels of significance across individuals, reflecting the diverse experiences of “Me Time”. These results underscore the role of personal environments like dorms in fostering meaningful routines and behaviors, offering insights into how individuals balance solitude, productivity, and emotional wellbeing.
@@ 5.2.2. Evaluation of Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Categories
Figure 8. Mean feature importance across all UIDs for each interaction label.
@@ 5.3. Scoring for Interaction Labels
To analyze how each feature contributes to scoring for interaction labels, we visualize the importance of relevant features across all 217 students. Figure 8 displays the top features (marked with a star) for each label:
@@
The scores for each interaction label are calculated based on the relative importance of various features, measured against a threshold value derived from the mean distribution of each feature across all UIDs. The scoring process increases the label’s score based on the contribution of features that meet the threshold condition, as outlined in Section 4.2.
@@
Figure 9 and Figure 10 depict the thresholds (represented as the mean values of the distributions, shown by the dashed red lines) for each feature within their respective interaction labels. In Figure 9, the dotted line highlights the threshold values for the most important features (F4, F13, F3, F5, F14, and F8). F4 reflects walking duration: if a student’s walking duration is less than the threshold (3 hours), it indicates leisure activity. Consequently, the LeisureScore is incremented proportionally to the feature importance of walking for the respective student. Similarly, F8 considers voice detection in others’ dormitories. If the duration exceeds the threshold (15 minutes), it suggests that the student is having leisure time with friends, causing an increase in LeisureScore by the value of importance of the corresponding feature. The same logic is applied to the other features, as shown in the figure and detailed further in Algorithm 1.
In Figure 10, the thresholds for the most important features (F6, F4, F7, and F1) are delineated by the dotted line. F6 indicates total sleep duration: if a student’s sleep exceeds 7 hours, representing the standard benchmark for healthy sleep, the SleepScore is incremented based on the student’s feature importance for sleep. F4, which measures the time spent in the room, shows that durations greater than 12.9 hours typically signify overnight dorm occupancy, often indicative of sleep rather than other activities. F7 corroborates healthy sleep with thresholds above 7.5 hours, while F1 reflects quietness, with values below 6 minutes suggesting an environment conducive to rest. Each scenario contributes to the SleepScore as outlined in Algorithm 2.

### v2

@@ 5.2. Interpretability
Figure 6. The individual-level heatmaps that highlight feature importance for each interaction label across 217 individuals (UIDs). The y-axis represents label-specific features, which vary across different interaction labels.
@@ 5.2.1. Evaluation of Stage 1: Input Features $\rightarrow$ Interaction Labels
  • Label: Leisure - Feature F4 (Duration of walking) emerged as the most important predictor for 90.36% of individuals, emphasizing walking as a key leisure activity. Other features, such as F13 (Phone usage at home), F3 (Duration of running), F5 (Duration of conversations), F8 (Voice detection at others’ dorms), and F14 (Phone usage at others’ dorms), are moderately important for 20–25% of individuals, indicating that leisure also encompasses social interactions and phone use. In contrast, features like F1 (Biking duration), F2 (Footsteps), F6 (Duration of calls), F9 (Conversations at others’ dorms), F10 (Audio detections at social places), F11 (Duration at “leisure”), and F12 (Workout duration) consistently show lower relevance. F11 has low importance despite its direct link to the label, likely due to unclear dataset definitions, underscoring the need for improved data clarity. These findings illustrate how behaviors like walking and social interactions influence leisure and provide insights into broader themes related to mental health and lifestyle habits.
  • Label: Me Time - Features such as F16 (Phone usage at own dorm), F12 (Time spent at own dorm), F7 (Voice detections at own dorm), F13 (Study duration), and F17 (Phone usage at study locations) are important for 89.5% of individuals, with F12 and F16 being the most significant. These findings suggest that activities in dorm spaces, such as studying and socializing, are key indicators of “Me Time” and contribute to a supportive environment. Conversely, features like F1 (Biking duration), F2 (Footsteps), F3 (Running duration), F5 (Walking duration), and F14 (Workout duration) show low importance, indicating that individual physical activities are less relevant to this label. Other features display varying levels of significance across individuals, reflecting the diverse experiences of “Me Time”. These results underscore the role of personal environments like dorms in fostering meaningful routines and behaviors, offering insights into how individuals balance solitude, productivity, and emotional wellbeing.
@@ 5.2.2. Evaluation of Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Categories
Figure 8. Mean feature importance across all individuals for each interaction label.
@@ 5.3. Scoring for Interaction Labels
To analyze how each feature contributes to scoring for interaction labels, we visualize the importance of relevant features across all 217 students. Figure 8 displays the top features (marked with a star) for each interaction label:
@@
As illustrated in Algorithm 2, the scores for each interaction label are calculated based on the relative importance of various features, measured against a threshold value derived from the mean of distribution across all individuals. Next, we show the thresholds for Leisure Time and Sleep to demonstrate how feature contributions influence the final label scoring.
@@
In Figure 9, the dotted lines highlight the threshold values for the most important features (F4, F13, F3, F5, F14, and F8) for Leisure Time . F4 reflects walking duration: if a student’s walking duration is less than the threshold (3 hours), it indicates leisure activity. Consequently, the LeisureScore is incremented proportionally to the feature importance of walking for the respective student. Similarly, F8 considers voice detection in others’ dormitories. If the duration exceeds the threshold (15 minutes), it suggests that the student is having leisure time with friends, causing an increase in LeisureScore by the value of importance of the corresponding feature.
In Figure 10, the dotted lines highlight the threshold values for the most important features (F6, F4, F7, and F1) for Sleep. F6 indicates total sleep duration: if a student’s sleep exceeds 7 hours, representing the standard benchmark for healthy sleep, the SleepScore is incremented based on the student’s feature importance for sleep. F4, which measures the time spent in the room, shows that durations greater than 12.9 hours typically signify overnight dorm occupancy, often indicative of sleep rather than other activities. F7 corroborates healthy sleep with thresholds above 7.5 hours, while F1 reflects quietness, with values below 6 minutes suggesting an environment conducive to rest.